OpenStudy (anonymous):

Edward received a gift card for $85 to a movie theater. He may use the gift card for either admission tickets or snacks. Tickets cost $7.50 and his favorite snack costs $3.75. If he wants to use his gift card for at least 15 individual purchases, which is a combination of tickets and snacks Edward may purchase? 7 tickets and 8 snacks 10 tickets and 6 snacks 8 tickets and 6 snacks 11 tickets and 3 snacks

OpenStudy (jziggy):

You have two equations to set up for this t + s = 15 and 7.5t + 3.75s = 85 Next, I get s = 15 - t and substitute that for s in the second equation 7.5t + 3.75(15 - t) = 85 7.5t + 56.25 - 3.75t = 85 3.75t = 28.75 t = 7 and 2/3 Therefore, there is no exact answer and instead you can try trial and error. The first answer only costs 82.50 total and is as such a correct answer.

OpenStudy (anonymous):

so you would choose the first choice?

OpenStudy (jziggy):

Yes, I would

OpenStudy (anonymous):

It would be the first choice